OpenStudy (anonymous):

Show that the function sinx+cosx is of period 2pi. Also prove that sinx+cosx=sqrt(2)sin(x+pi/4)

OpenStudy (anonymous):

hi satellite help me after this

OpenStudy (goformit100):

i think it is C.B.S.E. Question

OpenStudy (anonymous):

\[\sqrt2\sin(x+\frac{ \pi }{ 4 })\]

OpenStudy (anonymous):

what does C.B.S.E means ?

OpenStudy (anonymous):

sinx cos pi/4 + cosx sin pi/4 what's next?

OpenStudy (anonymous):

it is a consequence of the "addition angle" formula the general case is \[a\sin(x)+b\cos(x)=\sqrt{a^2+b^2}\sin(x+\theta)\] and a succinct explanation is in the attachment

OpenStudy (campbell_st):

this requires putting into the form asin(x) + bcos(x) = Rsin(x + m) \[R = \sqrt{a^2 + b^2}\] and \[m = \tan^{-1}(\frac{a}{b})\] so\[R = \sqrt{1^2 + 1^2 } = \sqrt{2}\] \[m = \tan^{-1}(\frac{1}{1}) = \frac{\pi}{4}\] so you get \[\sqrt{2}\sin(x + \frac{\pi}{4})\]